Suppose you draw 2 cards from a deck of 52 cards.
a) What is the probability that both cards are spades?
b) What is the probability that at least one card is a spade?

Answer by stanbon(75887)

Suppose you draw 2 cards from a deck of 52 cards.
a) What is the probability that both cards are spades?
13C2/52C2 = 0.0588
----------------------------------
b) What is the probability that at least one card is a spade?
1-P(no card is space)
= 1-39C2/52C2
= 1 - 0.5588
= 0.4412
---------------
Cheers,
Stan H.